Solve the diffrential equation (2x- y + 1)dx + (2y - x - 1) dy = 0.

Grade:12th pass

1 Answers

Aditya Gupta
5 years ago
So we can write it as 2xdx+dx+2ydy-dy= xdy+ydx
But d(xy)= xdy+ydx
So 2xdx+dx+2ydy-dy= dxy
Integrate both sides
x^2+x+y^2-y= xy+C